OpenStudy (anonymous):

What is the center and radius of the circle with the equation x2 + y2 + 8x - 4y - 61 = 0?

OpenStudy (anonymous):

Use Completion Of Square Method here..

OpenStudy (ghazi):

just rearrange the equation and try to make it in (x-x1)+(y-y1)= r^2

OpenStudy (anonymous):

Missing squares !!!! Wow...

OpenStudy (ghazi):

oh sorry ....

OpenStudy (anonymous):

Ghazi means : Make the equation looks like: \[(x - h)^2 + (y-k)^2 = r^2\] Here: \((h, k)\) is center and \(r\) is radius..

OpenStudy (anonymous):

@Qawi when will you participate?? When we will get offline ??

OpenStudy (anonymous):

i got ( 4,2) radius = 9

OpenStudy (ghazi):

(x^2+8x+4)+(y^2-4y+2)- 6=61 ...i have rearranged and added four and two....so now you can do this...yes you're right

OpenStudy (ghazi):

but radius = sqrt 67

OpenStudy (anonymous):

thanks :-)

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

Center is not right..

OpenStudy (ghazi):

and you've forgotten one more thing here...that is sign of coordinates

OpenStudy (ghazi):

it's (-4,2)

OpenStudy (ghazi):

@waterineyes am i right?

OpenStudy (anonymous):

Yes but in radius @Qawi is right..

OpenStudy (anonymous):

Radius is 9...

OpenStudy (ghazi):

how radius is 9?

OpenStudy (anonymous):

let me tripple check

OpenStudy (anonymous):

There you will add or subtract \(4^2\) and not 4..

OpenStudy (anonymous):

srry dissconnected but as i was saying i dont think it would be -4 , 2 because that wlld be the equain for x2 + y2 - 8x - 4y - 61 = 0? We're looking for x2 + y2 + 8x - 4y - 61 = 0

OpenStudy (ghazi):

no i guess radius is sqrt 67

OpenStudy (amistre64):

+ 8x - 4y half those terms and negate them

OpenStudy (amistre64):

if you were looking for the equation of the circle in standard form ... then youd have to go all out; but the center parts are just the negation of half the "linear" terms for x and y

OpenStudy (anonymous):

(x+4)^2 - (x-2)^2?

OpenStudy (amistre64):

pfft, it says center AND radius ..... then yes, water was correct ;)

OpenStudy (anonymous):

lol so its (-4,2); Radius =9

OpenStudy (anonymous):

Yep..